OpenStudy (anonymous):

EASY QUESTION: how many μg are there in 1 g?

OpenStudy (unklerhaukus):

the si prefix µ (micro) mean 10^-6

OpenStudy (unklerhaukus):

a micro gram is much smaller than a gram so there should be lots of micrograms in a gram

OpenStudy (anonymous):

ok got it :)

OpenStudy (unklerhaukus):

whats the answer @Data_LG2 ?

OpenStudy (anonymous):

1/(1 x 10^-6) g ?

OpenStudy (unklerhaukus):

yes good, now can you convert that fraction to an integer

OpenStudy (anonymous):

1000000g i just need this conversion because i need to convert something :) anyways, uhm.. are you taking any physics course? @UnkleRhaukus

OpenStudy (unklerhaukus):

it's a million, like you had but the units cancel out and yes im studying nuclear science

OpenStudy (unklerhaukus):

1 g / 1µg = 1/µ =1/10^-6 = 10^6

OpenStudy (anonymous):

but i thought micro gram is smaller than gram so it should be 0.0000001 g =1 µg oh yeah, i'm looking for 1 g, not 1 µg oh man! i'm making stupid mistakes again, smh thanks!